The greatest ratio among the ratios 2 : 3, 5 : 8, 75 : 121 and 40 : 25 is
(A) 2 : 3 (B) 5 : 8 (C) 75 : 121 (D) 40 : 25

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (D) is correct here.
Consider the given ratios, 2 : 3, 5 : 8, 75 : 121 and 40 : 25.
Simplified form of 2: 3 = 2/3
= 0.67
Simplified form of 5 : 8 = 5/8
= 0.625
Simplified form of 75: 121 = 75/121
= 0.61
Simplified form of 40: 25 = 40/25
= 1.6
Therefore, the greatest ratio among the given ratios is 40 : 25


Thanks